OpenStudy (anonymous):

use separation of variables to solve the initial value problem. dy/dx = (y+2)(x+2) and y=1 when x=0

OpenStudy (anonymous):

*oops I meant (y+5)

OpenStudy (anonymous):

dy/dx= (y+5)(x+2) and y=1 when x=0

OpenStudy (anonymous):

dy/(y+5) = (x+2) dx integrate both sides

OpenStudy (anonymous):

I did, but I don't know what to do after.. I did this?